The initial allele (an alternate form of a gene) proportions in a genetic simulation follow a...

  1. The initial allele (an alternate form of a gene) proportions in a genetic simulation follow a beta distribution with α=4 and β=3.
  1. What is the probability that this proportion will exceed 60%?
  2. Find mean proportion.
  3. Find the most likely proportion (mode).
  4. Find the variance.    
Answer #1

a)

P(p > 0.6)

= 0.45568

In R

1-pbeta(0.6,4,3)
[1] 0.45568

b)mean = a/(a+b) = 4/(4+3) = 4/7 = 0.5714

c)

mode = (a-1)/(a+b-2) = 3/(4 + 3 - 2) = 3/5 = 0.6

d)

variance = ab/((a+b)^2 (a +b+1))

= 4 * 3 / ( ( 4+3)^2 * (4 + 3 +1))

= 0.03061